Calculate the standard enthalpy change, ΔH, for the following reaction:




2CH3OH(l) + 3O3(g) —> 2CO2(g) + 2H2O(g)




Given:



CH3OH (l) ΔHf° = -238.7 kJ



CO2 (g) ΔHf° = -393.5 kJ



H2O (l) ΔHf° = -285.8 kJ





1
Expert's answer
2022-05-06T12:27:02-0400

"2CH_3OH+3O_2 \rightarrow 2 CO_2+4H_2O"

"\Delta H=4\Delta H[H_2O]+ 2\Delta H [CO_2]- 3\Delta H[O_2]-2\Delta H [CH_3OH]=4(-285.8)+2(-393.5)-3(0)-2(-238.7)=-1143.2-787-0+477.2=-1453 kJ"